OpenStudy (anonymous):

Two forces, F⃗ 1 and F⃗ 2, act at a point. F⃗ 1 has a magnitude of 9.80N and is directed at an angle of α = 64.0∘ above the negative x axis in the second quadrant. F⃗ 2 has a magnitude of 5.00N and is directed at an angle of β = 53.4∘ below the negative x axis in the third quadrant. What is the x component Fx of the resultant force?

OpenStudy (anonymous):



OpenStudy (anonymous):

Remember the x component of each force is given by \[F_{1x}=F_1\cos \alpha \\ F_{2x} = F_2 \cos \beta\] and that the x component of the resultant force, F1x + F2x, points in the negative x direction.